OpenStudy (anonymous):

Angular and Linear Velocity... A small gear of radius 5 cm is turning with an angular velocity of 20 radians per second. It drives a large gear of radius 15 cm. What is the linear velocity of the teeth on the large gear?

OpenStudy (amistre64):

linear velocity is how fast your going on the outer edge

OpenStudy (amistre64):

we travel 2pi r, the circumference, in a given time span distance* time = speed

OpenStudy (amistre64):

the 2 gears should be linearly the same

OpenStudy (amistre64):

2pi/20 = pi/10 parts of the circumference per second, right?

OpenStudy (amistre64):

5*pi/10 = pi/2 per sec

OpenStudy (amistre64):

that almost makes sense to me lol ...

OpenStudy (amistre64):

2pi = 6. somehting so its moving quicker; i think i got my ration upside down

OpenStudy (anonymous):

I got a linear velocity for the large gear of 300 cm per second when I used the equation using angular velocity of the small gear... I think I'm doing it correctly but I'm really not sure?

OpenStudy (amistre64):

i just cant recall the formulas so i gotta reinvent it :)

OpenStudy (anonymous):

The formula given says v=w(r), v being the linear velocity, w being angular, and r being the radius. Trig is kicking my butt.

OpenStudy (amistre64):

6.3662 pi per second 2pi * n = 6.3662 pi n = 6.3662/2 = 3.1631 or thereabouts yeah; radius * angle swept out = distance travled

OpenStudy (amistre64):

20 is the angle in rads so I guess 20*5 = 100 rads per second

OpenStudy (amistre64):

both gears are traveling at the same linear speed; or else one would always be catching up to the other

OpenStudy (amistre64):

they differe in angular speeds simply becasue they rotate differently; but linearly they are equal

OpenStudy (amistre64):

they cover equal distances in the same aount of time on their edges

OpenStudy (anonymous):

Okay okay that makes a lot more sense now. But the angular speed changes because the distance the gear covers is different for different sizes? Is that correct?

OpenStudy (amistre64):

you ever see a small dog trying to keep up with a larger one?

OpenStudy (amistre64):

if they cover the same distance in the same amount of time they are linearly equal; but the smaller one has to move alot faster becasue of its size; they are angularly different

OpenStudy (anonymous):

That helps a lot!

OpenStudy (amistre64):

:) insanity has its benefits lol

OpenStudy (anonymous):

So would a drive and wheel sprocket and wheel work the same way? They are connected by a chain, so they would all have the same linear velocity as well..

OpenStudy (amistre64):

yes, same linear velocity; the smaller wheel just has to turn around quicker to cover the same distance so its angle speed is faster